# XL-Instruct: Synthetic Data for Cross-Lingual Open-Ended Generation

### v1

@@ 7 Conclusion
With this goal in mind, we make three key contributions. First, we introduce the XL-AlpacaEval benchmark to evaluate the current state of open LLMs, and report poor performances and significant gaps against GPT-4o-Mini. Second, we propose the XL-Instruct technique, and show that this synthetic data can substantially boost cross-lingual performance, both in terms of win rates and fine-grained quality metrics. Third, we show that it exhibits strong zero-shot transfer to monolingual generation, both in English and beyond. Based on these results, we strongly encourage researchers to post-train their multilingual LLMs with XL-Instruct data, and shall publicly release our versions of XL-Instruct and XL-AlpacaEval datasets to support this.
8 Ethics Statement
There has been some concern in the literature that iterative training on synthetic data could eventually lead to model collapse (Shumailov et al., 2024). Like any other synthetic data technique, XL-Instruct could also share similar risks, especially since its seed data is sourced from the Web. We conduct fine-grained evaluation along various quality metrics, including Objectivity and Naturalness, to ensure that our fine-tuned models continue producing neutral, unbiased, and natural-sounding outputs and observe performance comparable to a fully instruction-tuned model fine-tuned on both human and synthetic data – which helps alleviate such concerns. We note that later works have also shown that mixing synthetic with human-generated data could avoid model collapse (Seddik et al., 2024; Gerstgrasser et al., 2024). While outside the scope of this work, it would be interesting to explore this angle as part of future research, and observe how XL-Instruct interacts with data from other tasks
Appendix A The XL-AlpacaEval Benchmark
Here we provide some additional details on the creation of the XL-AlpacaEval benchmark, which has 797 cross-lingual prompts in total, and currently supports 11 languages - the 8 languages used for the primary experiments in this work (Chinese, German, Hindi, Hungarian, Irish, Lithuanian, Maltese and Portuguese) and 3 additional languages (French, Finnish and Turkish) which we use for zero-shot evaluation in future sections. It is trivial to extend it to other languages – one simply has to run a script to append cross-lingual generation instructions (Section A.2) to our filtered AlpacaEval test set (Section A.1) and such extensions are being planned as a part of future work.
A.1 Manual Verification
@@
200 Write ”Test”
350 I’m an English speaker trying to learn Japanese Kanji using mnemonics. Mnemonics for Kanji are created from the primitives that make them up. The Kanji for Tax has the primitives wheat and devil, so an example would be, ”Taxes are like the devil taking away your hard earned wheat”. Can you create a mnemonic for the Kanji meaning Wish that has the primitives clock and heart?
458 Give me a list of 5 words where the letters of the words are in alphabetical order. One example: ”doors”. ”d” comes before ”o”, ”o” comes before ”r”, and ”r” comes before ”s”.
@@
Before creating our cross-lingual benchmark, we conduct a rigorous stage of manual verification to ensure that the prompts are suitable for answering cross-lingually. In Table 4, we show the prompts we removed from AlpacaEval that were too culturally specific (for instance, prompt 183) or tailored towards eliciting an English response (prompts 350 and 714). In the latter, we felt mandating a non-English response might make evaluating a “correct” response challenging. In other cases where the prompt simply requested a response in English, we replaced with a generic templated variable {language} for downstream substitution with the name of the desired target language. This leaves us with a total of 797 prompts. It is important to note that as far as possible, we tried to keep complex multi-step, multilingual prompts in our evaluation set, and only removed cases that were clearly invalid – in keeping with the goal of this work to build robust cross-lingual models.
A.2 Generation prompts
Next, we randomly sample prompts from a list of cross-lingual generation instructions (given in Table 5), and append it to each prompt in the filtered test set from the previous stage. To add further diversity to the instructions in the benchmark, we remove the word “language” from the prompts given in Table 5 – thus converting “Answer in German language” to “Answer in German”. This leads to the creation of the final XL-AlpacaEval benchmark.
Appendix B Experiments
@@ B.1 XL-AlpacaEval Results
Model Avg French Finnish Turkish
EuroLLM 9B 7.80 9.69 9.78 3.94
EuroLLM 9B Instruct 14.08 19.39 14.05 8.81
EuroLLM 9B XL-Instruct (best) 20.62 25.8 22.72 13.33
Zero-Shot Results
We show in Table 7 evaluations on zero-shot performance after fine-tuning with XL-Instruct. We choose French, Finnish and Turkish, 3 languages the EuroLLM model is pre-trained on, and observe huge gains in win rates, largely outperforming even the EuroLLM 9B Instruct model. This shows that even if done only for a few languages, XL-Instruct can still result in significant transfer that improves performance in others. We hypothesize that this is likely because the model is able to learn formatting, response structure etc. from this process, which also supports the boosts in English generation one observes in Table 3.
Fine-grained Evaluation
Finally, we also tabulate the macro-averaged GPT-4o scorings on Precision, Informativeness, Naturalness and Objectivity metrics (evaluation prompts and rubrics available at https://tinyurl.com/xl-gen-eval). As noted previously, the EuroLLM 9B model performs the worst, but the XL-Instruct model performs comparably or slightly better than EuroLLM 9B Instruct, indicating the efficacy of our data generation method.
Model Average Precision Informativeness Naturalness Objectivity
EuroLLM 9B 2.43 2.52 2.69 2.25 2.27
EuroLLM 9B Instruct 3.56 3.68 3.80 3.54 3.23
EuroLLM 9B XL-Instruct (best) 3.60 3.63 3.88 3.64 3.24
Table 8: Performance of EuroLLM 9B models evaluated on Precision, Informativeness, Naturalness, and Objectivity, with the average of these metrics.
B.2 Comparison with X-Instruction
Model Avg Finnish Hindi Turkish
EuroLLM 9B + X-Instruction (full 1M) 9.73 14.35 7.22 7.63
EuroLLM 9B + X-Instruction (40K) 10.44 13.69 8.76 8.86
EuroLLM 9B + XL-Instruct (naive, 40K) 12.06 15.3 10.9 9.98
EuroLLM 9B + XL-Instruct (best of 3, 40K) 17.82 23.15 15.8 14.52
Table 9: Performances of EuroLLM 9B models fine-tuned on X-Instruction and XL-Instruct data
We also compare the efficacy of our method with the most similar work to ours and the only cross-lingual open-ended generation dataset we are aware of: X-Instruction (Li et al., 2024a). In this work, the authors prompt a teacher LLM for reverse instructions directly in the low-resource language. Given this initially results in instructions of much poorer quality due to poor teacher capabilites, they also conduct iterative scoring and refinement to improve quality, achieving impressive results. They also release their dataset publicly at https://huggingface.co/datasets/James-WYang/X-Instruction.
We use the Hindi, Finnish, and Turkish splits of this dataset since these are supported by EuroLLM and are also available in X-Instruction. We also generate XL-Instruct data in these languages, by redoing the XL-Instruct pipeline (Section 4) from Stage 3 (Response Translation) for these languages. We LoRA fine-tune EuroLLM 9B on various X-Instruction and XL-Instruct datasets. For the former, we use both the entire 1M sized dataset available for these languages (in total), and a 40K instructions subset which is more comparable to our XL-Instruct baselines. For XL-Instruct, we train two baselines – one trained on ‘naive’ translations (ie. using only EuroLLM 9B Instruct) and another using the ‘best-of-3’ method (please refer Section 4.3 for a detailed explanation).
We see that both XL-Instruct baselines significantly outperform X-Instruction, with our best model achieving a 70.68% improvement over the latter – showcasing the relative superiority of our pipeline. This also suggests it might be more effective to prompt a teacher model in English due to inherently superior capabilities, and we hypothesize it might allow for greater quality and diversity in responses, as well as allowing for more complex operations – like refinement following specifically defined, custom criteria.
B.3 Ablations
Lastly, we conduct an ablation to verify the importance of the translation selection strategy. Given the cross-lingual part of the dataset mainly comes from Machine Translations, and translations can be quite noisy, we experiment with 2 MT techniques, ‘naive’ and ‘best-of-3’ responses. We also include a ‘random’ sampling strategy, where random responses are chosen for subsampling, regardless of MT quality. We fine-tune the EuroLLM 9B and EuroLLM 9B Instruct models using 8K and 32 instructions respectively, which are respectively the optimal SFT data sizes for each model (check Figure 2).
For the instruct model, ‘best of 3’ introduces significant improvements over naive or random sampling strategies, taking the average win rate from 18.55 to 20.84. This is likely because at the tiny scale of 32 instructions, target response quality matters hugely and significantly impacts performance. For EuroLLM 9B, which is fine-tuned on 8K instructions, performance still improves for most languages with the best-of-3 technique. The only cases where it drops are for the least-resourced languages like Irish and Maltese, which makes the average score much lower. It is possible the CometKiwi model we use for Quality Estimation is not very well-suited for such low-resource languages. As a result, we hypothesize that best-of-3 might sometimes end up choosing a worse translation than the naive method – which uses EuroLLM, a model known to have strong MT capabilities for all these languages.

### v2

@@ 7 Conclusion
With this goal in mind, we make three key contributions. First, we introduce the XL-AlpacaEval benchmark to evaluate the current state of open LLMs, and report poor performances and significant gaps against GPT-4o-Mini. Second, we propose the XL-Instruct technique, and show that this synthetic data can substantially boost cross-lingual performance, both in terms of win rates and fine-grained quality metrics. Third, we show that it exhibits strong zero-shot transfer to monolingual generation, both in English and beyond. Based on these results, we strongly encourage researchers to post-train and evaluate their multilingual LLMs on our publicly released XL-Suite.
8 Limitations
There has been some concern in the literature that iterative training on synthetic data could eventually lead to model collapse (Shumailov et al., 2024). Like any other synthetic data technique, XL-Instruct could also share similar risks, especially since its seed data is sourced from the Web.
We also did not perform human evaluation on the synthesized data or the model-generated outputs due to cost and time considerations. This drawback may have been partially mitigated by the rubric-based LLM judgments.
LLM Usage Statement
AI assistants were used to aid the programming and writing process in this work. For coding, it was used to create helper functions for preprocessing, and resolve bugs. During writing, it was used to aid in constructing LaTeX tables, plot graphs, fix grammar, etc.
Contributions
We list author contributions loosely following the CRediT author statement. [^6] https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement
  • VI: Conceptualization (co-lead), Software (lead), Writing (lead), Methodology (lead), Formal analysis (lead), Data Curation (lead), Visualization (lead)
  • PC: Conceptualization (co-lead), Writing (supporting), Supervision (supporting)
  • RR: Writing (supporting), Supervision (supporting)
  • AB: Writing (supporting), Supervision (lead), Project Administration (lead), Funding Acquisition (lead)
Acknowledgments
This work has received funding from UK Research and Innovation under the UK government’s Horizon Europe funding guarantee [grant numbers 10039436 and 10052546]. The first author, VI, also received funding from the Apple Scholars in AI/ML PhD fellowship program. Finally, we thank EDINA team at the University of Edinburgh for their provision of OpenAI credits through the ELM API that facilitated all the experiments in this work.
We thank Wenhao Zhu for discussions during prior work (Iyer et al., 2024b) that helped in the formulation of this idea. We also thank Simran Khanuja for her feedback on initial drafts of this work.
Appendix A Data
A.1 The XL-AlpacaEval Benchmark
Here we provide some additional details on the creation of the XL-AlpacaEval benchmark, which has 797 cross-lingual prompts in total, and currently supports 11 languages - the 8 languages used for the primary experiments in this work (Chinese, German, Hindi, Hungarian, Irish, Lithuanian, Maltese and Portuguese) and 3 additional languages (French, Finnish and Turkish) which we use for zero-shot evaluation in future sections. It is trivial to extend it to other languages – one simply has to run a script to append cross-lingual generation instructions (Section A.1.2) to our filtered AlpacaEval test set (Section A.1.1) and such extensions are being planned as a part of future work.
A.1.1 Manual Verification
@@
200 Write "Test"
350 I’m an English speaker trying to learn Japanese Kanji using mnemonics. Mnemonics for Kanji are created from the primitives that make them up. The Kanji for Tax has the primitives wheat and devil, so an example would be, "Taxes are like the devil taking away your hard earned wheat". Can you create a mnemonic for the Kanji meaning Wish that has the primitives clock and heart?
458 Give me a list of 5 words where the letters of the words are in alphabetical order. One example: "doors". "d" comes before "o", "o" comes before "r", and "r" comes before "s".
@@
Before creating our cross-lingual benchmark, we conduct a rigorous stage of manual verification to ensure that the prompts are suitable for answering cross-lingually. In Table 7, we show the prompts we removed from AlpacaEval that were too culturally specific (for instance, prompt 183) or tailored towards eliciting an English response (prompts 350 and 714). In the latter, we felt mandating a non-English response might make evaluating a “correct" response challenging. In other cases where the prompt simply requested a response in English, we replaced with a generic templated variable {language} for downstream substitution with the name of the desired target language. This leaves us with a total of 797 prompts. It is important to note that as far as possible, we tried to keep complex multi-step, multilingual prompts in our evaluation set, and only removed cases that were clearly invalid – in keeping with the goal of this work to build robust cross-lingual models.
A.1.2 Generation prompts
Next, we randomly sample prompts from a list of cross-lingual generation instructions (given in Table 8), and append it to each prompt in the filtered test set from the previous stage. To add further diversity to the instructions in the benchmark, we remove the word “language" from the prompts given in Table 8 – thus converting “Answer in German language" to “Answer in German". This leads to the creation of the final XL-AlpacaEval benchmark.
A.2 License
The XL-AlpacaEval dataset, which is derived from the AlpacaEval dataset, is released under a CC-by-NC 4.0 license, following its predecessor. This means the dataset is primarily intended for use in non-commerical (research) contexts. In contrast, the XL-Instruct dataset, which is provided as a training dataset, is derived from the CulturaX corpus – which in turn sources from the mC4 (Xue et al., 2021) and OSCAR (Ortiz Suárez et al., 2020). mC4 is released under an ODC-BY license, and OSCAR is released under CC0 no rights reserved. Hence, XL-Instruct can be used in both commercial and research contexts, as long as the corresponding licenses are respected.
Appendix B Additional Experiments and Results
@@ B.1 XL-AlpacaEval Results
B.2 Ablations
Lastly, we conduct an ablation to verify the importance of the translation selection strategy. Given the cross-lingual part of the dataset mainly comes from Machine Translations, and translations can be quite noisy, we experiment with 2 MT techniques, “naive” and “best-of-3” responses. We also include a “random” sampling strategy, where random responses are chosen for subsampling, regardless of MT quality. We fine-tune the EuroLLM 9B and EuroLLM 9B Instruct models using 8K and 32 instructions respectively, which are respectively the optimal SFT data sizes for each model (check Figure 2).
For the instruct model, “best of 3” introduces significant improvements over naive or random sampling strategies, taking the average win rate from 18.55 to 20.84. This is likely because at the tiny scale of 32 instructions, target response quality matters hugely and significantly impacts performance. For EuroLLM 9B, which is fine-tuned on 8K instructions, performance still improves for most languages with the best-of-3 technique. The only cases where it drops are for the least-resourced languages like Irish and Maltese, which makes the average score much lower. It is possible the CometKiwi model we use for Quality Estimation is not very well-suited for such low-resource languages. As a result, we hypothesize that best-of-3 might sometimes end up choosing a worse translation than the naive method – which uses EuroLLM, a model known to have strong MT capabilities for all these languages.